Mid-Mission Science Report – February 4th

Title: Self-Sustainment Study

Crew Member: Sarah “Ceres” Guthrie

Role: Commander

This study focuses on the efficiency of a self-sustainment diet for astronauts in spaceflight and on interplanetary missions. Meals are prepacked, dehydrated, and curated to the individual astronaut’s needs. Understanding the metabolic needs of astronauts and requirements for self-sustainment will enable a low-impact resource management capability for early exploration. Meals use no grid energy and in some cases are heated purely by solar energy (placing canisters in the window). The analysis of this study is completed at the end of the mission. The prepacked meals have been ample in caloric intake and consumption has varied based on activity. Moreover, as noted by the crew, the self-sustainment study shows there is little to no prep time, little waste, or cleanup, and has no impact on the daily activities. While halfway through the mission, there has been no need to break from the meal plan and is on track to be sufficient until the end of mission. Moreover, concerns if the meal would be disruptive have proved to be just the opposite. Not only is it supported by the crew as a specialized diet, but it has also generated meaningful conversations and even accolades, with some crew members quoted “I wish I was doing this.”

Title: Dandelion Study

Crew Member: Sarah “Ceres” Guthrie

Role: Commander

This study was a community submission by a 2nd grader from Millersville, Maryland who aspires to be an astronaut. The submission inquired if dandelions (Taraxacum officinale) can grow in Martian regolith. Crew 274 was happy to comply and donated Martian and Lunar regolith from ExoLabs to the study. The dandelions were planted in a nitrogen enriched pods with the sample regolith simulant on Sol 3 and are patiently waiting for results.

Title: Cosmic Fruit Study

Crew Member: Sarah “Ceres” Guthrie

Role: Commander

This study was a community submission by a 2nd grader from Millersville, Maryland who aspires to go to Mars as an astronaut. The submission asked what happens to fruit exposed to the elements on Mars and if they would be protected in the habitat. Crew 274 was happy to support this study in conjunction with the logistics cache. HSO Pender accommodated the fruit (two oranges) in his logistics cache and placed the cache and the fruit on the ”Martian terrain.” Two additional oranges were placed in a crate next to the cache. Both samples had temperature monitors and radiation tags. As the EVA crews access the cache over the mission duration, the oranges will be monitored and data collected on their performance.

Title: Evaluating Contingency EVAs and Rescue Techniques for Planetary Surface Missions

Crew Member: Sarah “Ceres” Guthrie

Role: Commander

This study investigates contingency extra-vehicular activities methods and protocols. It is a continuation of a study from previous analogs but enhanced over time. This specific study will utilize an engineered rescue vest. It has yet to be performed at this point in the mission.

Title: Digital Measurement of Stress and a Potential Stress Mitigation Technique in Analog Astronaut Environments

Crew Member: Alexis “Kepler” Lojek

Role: Crew Engineer

The crew are wearing their Garmin devices and it is tracking their stress levels based upon Heart Rate Variability. Focused breathing as a potential; mitigation technique for stress begins this evening after the Comms windows closes. Stress levels are being tracked and recorded via Garmin Connect.

Title: Heliophysical Phenomena

Crew Member: Noah “Phoenix” Loy

Role: Crew Astronomer and Heliophysicist

Musk Solar Observatory operations have rendered many successes at this halfway point mid-mission. Dozens of heliophysic phenomena have been observed in the forms of solar dark spots, solar chromosphere granolas, solar prominences, and convection cells. Over 110k images/videos have been captured. Some of these have been analyzed through astrophotography software. Flats have been integrated across 30k solar dark spot images. The size, distribution, and frequencies of solar granules and solar dark spots have begun to be measured. All in all, lots of clean and useful heliophysics data have been gathered, saved, and analyzed for a solar cycle 25 report.

Title: Case Study of the MDRS Design as a Planetary Surface Habitat

Crew Member: William “Titan” O’Hara

I have made significant progress on my research goal. In the first half of the mission I have completed a detailed review of the architecture of the Musk Observatory, Science Dome, Green Hab, RAM and connecting tunnels from the POV of a crew member. In each case I have created sketches and completed a detailed questionnaire built to systematically review each habitable space. The data collected thus far captures characteristics such as layout, use-of-space, activity volume allocations, traffic flow, outfitting and stowage volumes.

Title: Generating Multi-bandpass Lightcurve (LC) Data on HADS Variable Star V0799 AUR

Crew Member: Salina “Nova” Peña

Role: Crew Astronomer

Images of HADS Variable Star V0799 AUR were taken before the mission. Fifteen images were taken on January 13, 2023, using the filter “V,” and sixty were taken on January 27, 2023, using filters B, V, and R. Once the correct duration for each filter was established, then 180 images were taken at MDRS using the same filters. So far, there have been two days where no observations were made due to weather conditions. Therefore, they were canceled. While in the waiting process, I started the calibration process with the first set of images in the 45s duration visible.

Title: Germination Studies of Long-Duration Space-Exposed Seeds

Crew Member: Tyler “Houston” Hines

Role: GreenHab Officer

Approaching the mid-way point of the mission, all expected research progress and general germination are going according to plan. Following the initial setup of supplies along with consistent daily maintenance and watering of nutrient-rich additives, early signs of germination were noted as early as Sol 3 in the mission, specifically with the cress and broccoli, with the other sets of micro greens being closely monitored in parallel. As a significant and optimistic milestone in the anticipated seed germination timeline, early signs of LDEF seed germination were noted on the morning of Sol 6, with the plan to continually maintain current temperature and humidity levels in addition to watering times to support future growth.

Title: Supply Cache Use for Extension of Human Exploration on Mars

Crew Member: Nicholas “XMan” Pender

Role: Health and Safety Officer

Three of seven EVAs planned for this research experiment have been accomplished. The first EVA series (EVAs 3 and 4 combined) established a baseline distance of how far is reasonably possible to travel on foot in 30 minutes while in a space suit. A distance of 1 mile per 30 minutes of walking was discovered, which allowed for future planning of the contingency scenario route utilizing the supply cache. This phase also presented the opportunity to test staking into the local soil for future supply cache solar panel placement. This test was conducted on top of both Kissing Camel Ridge and on the side of Cow Dung Road. Finally, I was also able to prove the ability to use a hydration system with the space suit and consume GU energy gel packs while in the space suit. EVA 5 demonstrated effective deployment of the supply cache as a mobile EVA support platform, positioning it 2 miles (1 hour) from the hab. Upon activation, all powered systems for the cache were operating nominally. Initial readings on interior/external temperature were performed, as well as power consumption data and initial condition. EVA 6 comprised of a status check on the cache to ensure it was operating nominally, and it certainly was. The cache was maintaining temps above 40 degrees and power consumption of the internal heater was negligible. One external concern, it was found that the ropes securing the solar panels to the ground had loosened. Stakes were repositioned to ensure a secure placement. Follow-up surveys on and debriefs on EVA’s 5 and 6 validated to the experiment design and also revealed potential design improvement recommendations for the cache system. The cache is now ready for its big test in an emergency scenario exercise on EVA 7!

Operations Report – February 4th

Crew 274 Operations Report 2/4/2023

SOL: 6

Name of person filing report: Alexis Lojek

Non-nominal systems: N/A

Notes on non-nominal systems: N/A


Spirit rover used: No

Hours: 208.0

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

Opportunity rover used: No

Hours: 112.2

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

Curiosity rover used: Yes

Hours: 220.3

Beginning charge: 100%

Ending charge: 55%

Currently charging: Yes

Perseverance rover used: Yes

Hours: 254.6

Beginning charge: 100%

Ending charge: 41%

Currently charging: Yes

General notes on rovers: Rovers Curiosity and Perseverance were used during today’s EVA. The team returned before noon, so they are plugged in and charging.

Summary of Hab operations: Habitat is in a nominal condition.

WATER USE: 33 gallons as of 6:00 pm.

Water (static tank): 314 gallons remaining

Static tank pipe heater (On or off): On

Static tank heater (On or off): On

Toilet tank emptied: Yes

Summary of internet: Nominal

Summary of suits and radios: Suits are in nominal condition and charging. Radios were charged after EVA completion and taken off the chargers after the red charging lights turned off. All conditions normal.

Summary of GreenHab operations: GHO Hines watered plants this morning from 1:30 – 2:00 pm.

WATER USE: 13.74 gallons used

Heater: On

Supplemental light: 4 hours (10pm-2am)

Harvest: Yes. Tomatoes – 232 g, green beans – 374 g, spinach – 19 g, micro greens – 21 g, basil – 13 g, dill – 1 g; total: 660 g.

Summary of ScienceDome operations: GHO Officer Tyler Hines’ experiment is continuing to be monitored; Cress micro greens have made significant progress and have begun to sprout.

Dual split: On a timer from 10pm to 7am and under Mission Support supervision.

Summary of RAM operations: No operations were conducted in the RAM.

Summary of any observatory issues: No observatory issues.

Summary of health and safety issues: N/A

Questions, concerns, and requests to Mission Support: N/A

Astronomy Report – February 4th

Astronomy Report

Name: Salina Peña and Noah Loy

Crew: 274
Date: 02/04/2023


Robotic Telescope Requested (choose one) MDRS-14 MDRS-WF

Objects Viewed: Star/230204/Variable star

Problems Encountered: None (I canceled pictures tonight due to overcast weather. I am continuing the calibration process of images.)


Solar Features Observed: Sun/230204/sunspot Sun/230204/prominence

Problems Encountered: None (I will attach images later)

GreenHab Report – February 4th

GreenHab Officer: Tyler Hines

Environmental control: heater

Average temperatures: 82.2 F

Hours of supplemental light: 4 hours

Daily water usage for crops: 13.74 gallons

Daily water usage for research and/or other purposes: 1.06 gallons

Water in Blue Tank: 113.79 gallons

Time(s) of watering for crops: 1:30 pm

Changes to crops: N/A

Narrative: Completed daily watering of plants and crops as part of daily GreenHab operations. A significant mid-mission harvest was conducted as part of general GreenHab maintenance and to further support crew nutrition over the coming days.

Harvest: <1 gram of dill, 232 grams of tomato, 374 grams of green beans, 19 grams of spinach, 21 grams of micro greens, and 13 grams of basil.

Support/supplies needed: N/A

Journalist Report – February 4th

Crew 274 Journalist Report 2-4-2023

Sol: 6

Author: Tony DiBernardo, Crew Journalist

Title: Radiation-Exposed Seeds Germinated in Simulated Martian Regolith

Crew 274 woke up this morning at 7:30am to the song “Feel Good Inc.” by the Gorillaz and a recorded message from the family of the Crew Astronomer, Salina “Nova” Pena. Her daughter also sent some inspiring words, encouraging her mother to think of her time away as an opportunity to see home with new eyes upon arrival.

EVA Activity

Today HSO Pender led EVA 6 to collect temperature and power consumption data on EEVAC-1, the supply cache experiment intended to extend human exploration on Mars. After leaving the cache unattended for 24 hours from yesterday’s EVA, the cache was found to have performed as designed, though a debriefing of the crew following the EVA revealed potential design improvements on the current iteration of EEVAC-1. His EVA crew also collected temperature and radiation exposure data for Commander Guthrie’s community experiment put forth by a 2nd grade student from Millersville, MD. The rest of HSO Pender’s day was spent processing data collected from EVA 6 and preparing for tomorrow’s 3-hour walking EVA, which will be a climax study for the supply cache experiment on this mission.


Upon conducting daily morning observations, the first official evidence of germination of a single seed was noted of the LDEF seeds planted. While the evidence of germination was minor, this proves to be a significant step toward the sustainability of durability of nutrient-dense seeds exposed to high levels of radiation within the space environment. Add this to the fact that simulated lunar and martian regolith is practically “dead” or extremely hard to yield germination, this was an exciting and potential ground-breaking discovery. Following additional crew guidance, temperature and humidity levels will continue to be closely maintained to support further growth and germination of the other LDEF seeds. Additionally, the micro green cress continue to flourish in their development and are predicted to be ready for potential harvesting before the end of the mission. Meanwhile, the broccoli, red cabbage, and arugula are continually being tended to and monitored closely in their general progress.


Crew astronomers Loy and Pena observed detailed heliophysics phenomena in which multiple observations, videos, and photos were taken of solar dark spots, solar chromosphere granules, and solar prominences. Over 8 large solar prominences were recorded today, the largest dataset of prominences since experimentations began. These observations will be very useful for the astronomer’s analysis of solar cycle 25. Also of important note, Loy and Pena observed what seems to be an asteroid or a satellite crossing the line of sight between the solar observatory and the sun. This object is still being analyzed.


Cloudy weather conditions canceled all planned observations of HADS Variable Star V0799 AUR. Crew Astronomers instead continued their analysis of their current image inventory, stacking and analyzing 9 separate 45-second images in the visual filter. So far, HADS Variable Star V0799 AUR is showing signs of dimming, which provides good data to the astronomers’ variable star report.

Hab Design

Hab Specialist Bill O’Hara participated in EVA 6 in support of HSO Pender’s research. During this EVA, he took sensor measurements from the cache and drove the rover back to the hab. In the afternoon, Bill continued his case study of the MDRS architecture by reviewing the layout of the lower level of the Habitat. This level of the habitat module includes the airlock, suit stowage and the only bathroom used by the crew.

Stress Mitigation Study

Lex “Kepler” Lojek also participated in EVA 6, driving one way to the cache, assisted in taking measurements of the cache, and then helped navigate back. After returning, he continued to collect and organize stress data for his ongoing focused breathing study. Tonight, he’ll begin a daily study with the crew where they’ll conduct a 15-minute focused breathing meditation to study its effects on potential stress reduction during an analog mission.

EVA Report – February 4th

Crew 274 EVA Report 04-02-2023

EVA # 6

Author: Sarah E. Guthrie, Commander

Purpose of EVA: EEVAC-1 (supply cache) data collection to ensure safe usage for future EVAs.

Start time: 10:00am

End time: 11:30am

Narrative: EVA #5 purpose was to deploy the logistics cache in a specific location which will be retrieved in EVA #7. The crew exited the habitat at 10:00 and proceeded on their route to the cache via Rovers Percy and Curiosity at 10:04am. They arrived at the cache at 10:13am and began taking measurements of the cache which included collecting data readings on internal and external cache temperature, cache power consumption, internal and external cache condition, and temperature and radiation readings on the fruit experiment both inside and outside of the cache. The crew did have to retake into the ground the solar panels which the cord holding them down came loose, however the solar panels were undamaged. The crew proceeded back to the habitat at 11:13am and returned to the habitat at 11:23am, ingressing at 11:27am.

Destination: South on Cow Dung Road, near Zubrin’s Head.

Coordinates (use UTM WGS 84): N4248400, E519300

Participants: Nicholas Pender (HSO), Bill O’Hara (Journalist), Lex Lojek (Engineer), and Tyler Hines (Greenhab)

Road(s) and routes per MDRS Map: EVA #5’s (map provided) purpose is to drive to EEVAC-1 deployed at N4248400, E519300 via south on Cow Dung Road near Zubrin’s Head.

Mode of travel: Rovers Curiosity/Percy

Sol Summary Report – February 4th

Crew 274 Sol Summary Report 02-04-2023

Sol: 6

Summary Title: Cache Assessment, Green Beans, and Alien Spaceships?

Author’s name: Sarah E. Guthrie (“Ceres”), Commander

Mission Status: Nominal

Sol Activity Summary: The crew completed EVA #6 which was an assessment of the logistics cache performance overnight. The cache box (EEVAC-1) was able to maintain power and temperature successfully. The crew also took measurements of the oranges which was a community submission from a 2nd grader to explore the effects of fruit on Mars. The fruit seems unphased by the Martian elements through the protection of the cache. The crew enjoyed a “feast” of greenhouse harvest at dinner which brought a pleasant tone to the meal with fresh green beans (approx. 67) and cherry tomatoes (29). Tonight is the first day of the mediation and focused breathing study, which will be after reports are filed. The astronomy team captured several massive prominences, sun spots, granules, and solar flares. While reviewing the footage, the crew noticed a small speck flying across the Sun’s image. While it is not known yet what the fast moving object was, it was exciting and captured the crew’s imagination. Aliens?

Look Ahead Plan: The crew looks forward to the emergency scenario exercise of the cache system (pending approval).

Anomalies in work: None

Weather: Sunny, 1C High, -5C Low; Sunny, evening clouds

Crew Physical Status: Excellent

EVA: EVA #6 completed for logistics cache assessment – successful

Reports to be filed: Sol Summary, Operations, EVA Request, EVA Report, GHO Report, and Astronomy Report.

Support Requested: None

EVA Report – February 3th

Crew 274 EVA Report 03-02-2023

EVA # 5

Author: Sarah E. Guthrie, Commander

Purpose of EVA: Logistics Cache Deployment

Start time: 10:00am

End time: 11:30am

Narrative: EVA #5 purpose was to deploy the logistics cache in a specific location which will be retrieved in EVA #7.

Destination: South on Cow Dung Road, near Zubrin’s Head (see map provided).

Coordinates (use UTM WGS 84): N4248400, E519300

Participants: Nicholas Pender (HSO), Tony DiBernardo (Journalist), Noah Loy (Heliophysics), and Salina Pena (Astronomer)

Road(s) and routes per MDRS Map: EVA #5’s (map provided) purpose is to drive the cache and establish the power system at the way-point for EVA #7.

Mode of travel: Rovers Curiosity/Spirit *NOTE* Original Request was for Curiosity and Percy, the crew accidentally took the wrong rover.


Copyright © The Mars Society. All rights reserved. | Main Site